A hydrogen lamp emits several lines in the visible region of the electromagnetic spectrum. One of these lines has a wavelength of 6.77 x 10-7 m. What is the frequency of this radiation?
1
Expert's answer
2020-10-21T13:21:26-0400

Given the wavelength of electromagnetic line emitted by Hydrogen lamp, λ = 6.77 x 10^-7 m


As we know if frequency of electromagnetic radiation = ν s^-1

And velocity of it = C m/s


Then, C = v × λ

Or, v = C/λ


As, velocity of light, C = 3 × 10^8 m/s


So,

v = ( 3 × 10^8 m/s ) / ( 6.77 x 10^-7 m )

=4.431 × 10^14 s^-1

Hence the frequency of the given line

= 4.431 × 10^14 s^-1